Annual income in thousands: 12,13,14,15,16,17,18,19,20.
Annual savings in thousands:
0,0.1,0.2,0.2,0.5,0.5,0.6,0.7,0.8,
Find the following
1) Equation line y on x
2) Equation line x on y
3) Find the Annual saving when Annual income is 25000.
4) Find the Annual income when Annual saving is 550.
1
Expert's answer
2020-05-04T18:02:56-0400

Let "x=" annual income in thousands, "y=" annual savings in thousands.


"\begin{matrix}
 & x & y & xy & x^2 & y^2 \\
 & 12 & 0 & 0 & 144 & 0 \\
 & 13 & 0.1 & 1.3 & 169 & 0.01 \\
 & 14 & 0.2 & 2.8 & 196 & 0.04 \\
 & 15 & 0.2 & 3 & 225 & 0.04 \\
 & 16 & 0.5 & 8 & 256 & 0.25 \\
 & 17 & 0.5 & 8.5 & 289 & 0.25 \\
 & 18 & 0.6 & 10.8 & 324 & 0.36 \\
 & 19 & 0.7 & 13.3 & 361 & 0.49 \\
 & 20 &0.8 & 16 & 400 & 0.64 \\ 
 Sum= & 144 & 3.6 & 63.7 & 2364 & 2.08
\end{matrix}"

"mean: \bar{x}={\sum x_i\over n}={144\over 9}=16, \bar{y}={\sum y_i\over n}={3.6\over 9}=0.4"

"SS_{xx}=\displaystyle\sum_{i=1}^nx_i^2-{1\over n}\big(\displaystyle\sum_{i=1}^nx_i\big)^2=""=2364-{(144)^2\over 9}=60"

"SS_{yy}=\displaystyle\sum_{i=1}^ny_i^2-{1\over n}\big(\displaystyle\sum_{i=1}^ny_i\big)^2=""=2.08-{(3.6)^2\over 9}=0.64"

"SS_{xy}=\displaystyle\sum_{i=1}^nx_iy_i-{1\over n}\big(\displaystyle\sum_{i=1}^nx_i\big)\big(\displaystyle\sum_{i=1}^ny_i\big)=""=63.7-{144(3.6)\over 9}=6.1"

1) Find the equaition line y on x 


"B={SS_{xy} \over SS_{xx}}={6.1\over 60}\approx0.101667"

"A=\bar{y}-B\cdot\bar{x}=0.4-({6.1 \over 60})(16)\approx-1.226667"

"y=-1.226667+0.101667x"

2) Find the equaition line x on y 


"M={SS_{xy} \over SS_{yy}}={6.1 \over 0.64}=9.53125"

"N=\bar{x}-M\cdot\bar{y}=16-({6.1\over 0.64})(0.4)=12.1875"

"x=12.1875+9.53125y"

3) Find the Annual saving when Annual income is 25000. 


"y=-1.226667+0.101667(25)=1.315008"

The Annual saving is "1315."


4) Find the Annual income when Annual saving is 550.


"x=12.1875+9.53125(0.550)=17.4296875"

The Annual income is "17430."